### v2

# Segmentation and Smoothing Affect Explanation Quality More Than the Choice of Perturbation-based XAI Method for Image Explanations

Abstract
Perturbation-based post-hoc image explanation methods are commonly used to explain image prediction models. These methods perturb parts of the input to measure how those parts affect the output. Since the methods only require the input and output they can be applied to any model, making them a popular choice to explain black-box models. While many different models exist and have been compared with one another, it remains poorly understood which parameters of the different methods are responsible for their varying performance.
This work uses the Randomized Input Sampling for Explanations (RISE) method as a baseline to evaluate many combinations of mask sampling, segmentation techniques, smoothing, attribution calculation, and per-segment or per-pixel attribution, using a proxy metric. The results show that attribution calculation, which is frequently the focus of other works, has little impact on the results. Conversely, segmentation and per-pixel attribution, rarely examined parameters, have a significant impact.
The implementation of and data gathered in this work are available online  [^1] https://github.com/guspih/post-hoc-image-perturbation [^2] https://bit.ly/smooth-mask-perturbation.
I Introduction
Over the past decade, deep neural networks (DNN) have proven proficient at solving computer vision tasks [1]. However, the black-box nature of DNNs causes issues, including difficulties in understanding when the model is wrong, lack of trust in the models, and legal issues [2]. The goal of the field of Explainable Artificial Intelligence (XAI) is to make AI models more transparent to mitigate these issues.
The complexity of DNNs makes it difficult to explain the model in its entirety. Instead, methods for explaining DNNs tend to focus on explaining individual predictions made by the model [3]. Some research focuses on developing inherently interpretable models [4]. However, these methods are bound to specific models and cannot be applied to the vast amount of non-interpretable models that exist and are being developed [3]. In these cases, it is common to use post-hoc explanations that are separate from the models and can be applied during prediction [5].
Many different types of post-hoc methods exist to explain computer vision models. Though many methods can only be used on models that satisfy certain criteria, such as differentiability in the case of gradient-based methods. This work instead concerns a type of post-hoc explanation method that does not place any restriction on the models being used; so-called perturbation-based methods. Perturbation-based explanations work by analyzing how the model’s predictions change as the original input is perturbed. For image prediction, this is commonly used to find the regions of the image that cause the model to make a certain prediction. This can be used to assess if a specific model or prediction is sound and help locate important features for a user.
Since the information in images is generally found in the relationships between many pixels [6], perturbing individual pixels is unlikely to have much impact on the prediction. As such, perturbations are typically made on larger pixel segments. Depending on the method these segments are either perturbed one at a time or several at once with different sampling methods for determining what segments to perturb.
Figure 1: The pipeline for perturbation-based image attribution used in this work. The image is segmented, samples indicating what segments to perturb are drawn, the sampled segments are perturbed, the model to explain makes predictions for the perturbed samples, and the input-output pairs are used to calculate per-segment and per-pixel attribution.
The general pipeline for calculating perturbation-based image explanations consists of segmenting, sampling, perturbing, predicting, and attributing, as shown in Fig. 1. The image is split into segments and a number of samples are drawn indicating which segments should be perturbed. For each sample, a new image is created by perturbing the indicated segments in some way. Perturbation often consists of occluding the segments with a solid color [7], but other distortions such as inpainting have also been used [8]. The model output from these perturbed inputs can then be used to attribute influence to the segments based on how the output changes when they are perturbed or not. There are many ways to calculate attribution based on the input-output pairs, such as average output when a segment is included [9] or excluded [10]. Another method is to train a surrogate model to predict the output based on the perturbations and use the learned parameters as attribution [11, 12].
The influence attributed to each segment is often visualized using attribution maps. Attribution maps are heat maps showing high influence using one color and fading to another color to show low influence. By analyzing the attribution map a human can understand the model’s behavior. For example, a prediction with a high influence on a non-salient part of the image can indicate problems with the model. Using attribution map explanations increases the joint human-AI performance on some tasks [13].
Since influence is attributed based on which segments are perturbed, most perturbation-based methods assign attribution to entire segments, but cannot differentiate the influence between pixels within a segment. The Randomized Input Sampling for Explanations (RISE) method solves this by using smooth perturbations, where pixels closer to the segment center are perturbed more [9]. With this approach, a per-pixel attribution can be calculated by assigning each pixel the influence of each segment weighted by how perturbed the pixel was for that segment.
As for most papers on perturbation-based explanations for images, in [9] RISE is described as an entire pipeline from segmenting to pixel attribution. As such it is not clear which parts of the RISE method, such as per-pixel attribution, that is responsible for its performance. This work explores the individual parameters of the perturbation-based pipeline to evaluate which parameters have the most impact on performance. This is done by taking a version of the original RISE pipeline and evaluating the original parameters alongside a variety of additional segmentation, sampling, perturbation, and attribution methods.
Due to the large number of parameter combinations to evaluate, human evaluation is not feasible for this study. Instead, this work uses a faithfulness metric dubbed Symmetric Relevance Gain (SRG) [8]. The evaluations are carried out on the ImageNet validation set [14] for three different CNNs [15, 16, 17] using both per-segment and per-pixel attributions. The different pipeline parameters that have been combined and evaluated are shown in Table I.
The results show that the method of calculating the attribution, which is typically what is highlighted as the most important part, has little impact on performance. Instead, segmentation and sampling, which are often overlooked, have a significant impact on performance. This calls into question studies that compare the original implementations of attribution calculation methods without correcting for the other parameters of those implementations. Additionally, per-pixel attribution gives a significantly higher average performance compared to per-segment attribution.
I-A Contributions
The primary contributions of this work are as follows.
  • Evaluating the different parameters of the perturbation-based image attribution pipeline finding that all parameters can have a significant impact on performance.
  • Adapting the smooth-mask enabled per-pixel attribution used in RISE to work with other perturbation-based attribution methods.
  • Improving the RISE pipeline by using Gaussian filtering instead of bilinear upsampling, which enables the use of many other segmenting methods.
  • Generalizing the Probabilistic Difference Analysis (PDA) [10] method to work with multiple perturbing multiple segments per sample.
TABLE I: The different parameters of the perturbation-based image explanation pipeline used in this work.
Segmenting + Perturbing Sampling Samples Model Attribution
Grid+Bilinear Only one 4000/8000 AlexNet CIU
Grid+Gaussian All but one 400 VGG-16 PDA
SLIC+Gaussian Random 50 ResNet LIME
Entropic SHAP
RISE
II Methodology
This work evaluates different segmenting, sampling, perturbing, and attribution methods using three different sample sizes as listed in Table I. All possible pipelines combining one of each parameter are evaluated by explaining the predictions made by three different ImageNet [14] pretrained CNNs as measured by the SRG metric [8]. The different parts of the experiments are described in detail in the following subsections.
II-A Segmenting
This work evaluates two segmenting techniques; grid and SLIC [18] segmentation. Grid segmentation splits the image a given number of times horizontally and vertically. SLIC is a rule-based algorithm that iteratively calculates segment "centers", assigns each pixel to the closest center in a color-position space, and recalculates the segment centers repeatedly until convergence.
The experiments use the same $7\times 7$ segmentation grid as the original RISE implementation [9]. To make the SLIC segmentation as similar to the grid implementation as possible, SLIC is instantiated with 49 segment centers in the experiments. The default scikit-image implementation for SLIC is used [19].
II-B Sampling
This work generates samples indicating which segments to perturb using “only one”, “all but one”, random, and entropic sampling. “Only one” sampling generates all samples where only one segment is perturbed. “All but one” sampling generates all samples where all but one segment is perturbed. Both “Only one” and “All but one” sampling also generate the sample where no segment is perturbed as a reference value. Random sampling generates a given number of samples where each segment is perturbed with a probability of $0.5$. Entropic sampling generates a given number of samples in increasing order of entropy, i.e., samples with as many or as few segments perturbed as possible.
Entropic sampling is implemented to be similar to the default KernelSHAP sampling behavior [12]. No segments are perturbed in the first sample, all segments are perturbed in the second, followed by all possible combinations of one segment perturbing and all combinations of one segment unperturbed, followed by combinations of two segments perturbed/unperturbed, and so on until the sample size is reached.
“Only one” and “all but one” sampling create one sample per segment, which makes them computationally efficient. They are presumably the simplest sampling methods that can be used and have worked well with Contextual Importance and Utility (CIU) method for detecting bleeding in gastro-enterological [20]. Random and entropic sampling can generate samples consisting of combinations of segments, which can yield many more unique samples. Using combinations of segments presumably allows for taking into account inter-dependencies between different segments, potentially leading to more informative explanations.
II-B1 Sample size
Random and entropic sampling use sample sizes of 4000/8000, 400, as well as 50. The 4000/8000 sample size is used to be consistent with the original RISE evaluation. AlexNet and VGG-16 use 4000 samples and ResNet models were evaluated with 8000 samples. “Only one” and “all but one” sampling always generate samples equal to the number of segments plus one which is 50 samples at most in this work.
II-C Perturbing
This work perturbs images by replacing the perturbed segments with a solid color, which is the most common way of perturbing images. The color used for perturbing is the mean color of the ImageNet training set, which is the color that will be normalized to 0 in the preprocessing of the Torchvision models used in this work, which is the same as the original RISE implementation. Like the original RISE implementation, all pixels in the segments are not replaced with exactly the given color and instead are faded between their original color and the normalization mean according to a segment mask that fades closer to the segment border. This mask is also later reused to attribute per-pixel in addition to the normal per-segment attribution.
Perturbing consists of pixel-wise multiplication between the normalized image and a perturbation mask of values between 0 and 1. The mask is created by setting all values in the segments to be perturbed to 0 and all others to 1, then the mask is smoothened so that the values closer to the center of each segment are close to 0 and those at the edges and beyond are closer to 1. Thus pixels outside the perturbed segments are mostly unchanged, but fade towards the normalization mean as they get closer to the segment centers. The original implementation achieves this by using bilinear upsampling to scale a $7\times 7$ grid of 0s and 1s to the size of the full image, an implementation that is replicated in this work. An issue with this method is that it relies on having a lower resolution mask to upscale which excludes using some popular segmentation methods such as SLIC. To combat this issue, this work introduces the use of Gaussian filters for mask smoothing. The Gaussian filter smoothens the mask by assigning each pixel in the mask the average value of all other pixels, weighted by a 2D Gaussian function centered on the pixel. This work uses a Gaussian filter with $\sigma=10$ which in this use case gives similar smooth masks when compared to bilinear upscaling.
II-D Attributing
This work evaluates five existing attribution methods, CIU [21], PDA [10], LIME [11], SHAP [12], and RISE [9]. In this work, the method names refer only to the actual attribution calculation of each method.
[…]
The Prediction Difference Analysis (PDA) method [10] works similarly to CIU, but uses average difference instead of maximum difference. PDA has been adapted to work with images [24], though both in the original and image implementation only a single feature is changed at a time (“only one” sampling). In this work, PDA has been generalized to work when multiple features are perturbed simultaneously. The PDA attribution is given by $w_{PDA}(s)=Y-avg(Y\backslash_{s})$.
Locally-Interpretable Model-agnostic Explanations (LIME) [11] was originally introduced as an umbrella term used to cover any instance where a single prediction is explained by training an interpretable model to mimic the original model’s prediction in the neighborhood of the original input. However, LIME has since been associated with specifically training a linear surrogate model [7, 4] as this is how the method was demonstrated originally. In this work, LIME is implemented as a linear model $y=b+\sum_{s\in S}w_{s}\cdot x_{s}$, where $y$ is the output of the model, $b$ and $w_{s}$ are the learned bias and weights, and $x_{s}=0$ if the segment is perturbed and $1$ otherwise. The attribution of LIME for segment $s$ is the value of $w_{s}$ after the linear model has been fit to the input-output pairs using least squares.
Kernel SHAP [12] is a modification to LIME such that, under certain assumptions, the weights learned by the linear model will tend towards the Shapley values [25] scoring how the features contribute to the prediction. This is achieved by scaling the input-output pairs with a kernel function $\pi(X)=\frac{|S|-1}{\binom{|S|}{|X|}|X|(|S|-|X|)}$, where $|s|$ is the number of segments and $|X|=\sum_{x_{s}\in X}x_{s}$. As such the SHAP values can be retrieved by solving $\pi(X)y=b+\sum_{s\in S}w_{s}\cdot\pi(X)x_{s}$ using least squares.
The attribution used by RISE [9] is similar to PDA, but instead of using the average decrease when the feature is perturbed, it uses the average prediction when it is not perturbed. This similarity means that in the case where every possible sample is used, the two methods have equivalent ranks of the segments by influence, which in turn produces equivalent SRG scores. RISE attribution for a segment is given by $w_{RISE}(s)=avg(Y\backslash_{\bar{s}})$.
In addition to per-segment ($w_{s}$) attribution, this work also evaluates attribution per-pixel similar to the original RISE implementation. Typically, perturbation-based explanations assign the same attribution to all pixels in the same segment ($w_{s}$). However, by using the smoothing masks ($M_{s}$) used for perturbation the attribution scores of a pixel can instead be set to the average of the attribution of each segment weighted by how perturbed the pixel was by that segment. This is formalized as $w_{p}=\frac{1}{\sum_{s\in S}M_{s}^{p}}\sum_{s\in S}w_{s}\cdot M_{s}^{p}$, where $M_{s}^{p}$ is the value of pixel $p$ in the perturbation mask of segment $s$. Note that this calculation means that pixels outside segment $s$ which were slightly perturbed due to the smooth mask, also include that influence in the calculation. This results in pixels at segment centers getting almost all their attribution score from that segment, while pixels at the borders of segments get a lesser attribution from all the segments they border.
II-E Evaluating
The various pipelines are tested by explaining the predictions of three ImageNet pretrained CNNs on the ImageNet validation set [14] and evaluating those explanations using the Symmetric Relevance Gain (SRG) metric [8]. The three pretrained CNNs are AlexNet [15, 26], VGG-16 [16], and ResNet-50 [17] using trained parameters from the Torchvision $0.15.2$ framework [27]. The input to the models is normalized using the average pixel values of ImageNet.
The ImageNet validation set is used as it is the same dataset on which RISE was originally evaluated. Evaluation is carried out using the first image of each class of the ImageNet validation for a total of 1000 images ($2\%$ of the total validation set). The limitation to 1000 images is to make the computation demands feasible as over 300 pipelines were evaluated, some requiring 8000 model calls per image. Despite the limited number of images and some pipelines using the same model calls as they share the first parameters, the evaluation still required close to a billion model calls. Limited evaluation was performed on the full validation set and no statistically significant ($p<0.05$) difference could be found compared to using 2% of the data. For each image, the top predicted class of each model was explained through segment and pixel attribution using each pipeline.
The SRG metric is used due to the large number of parameter combinations to evaluate, which signifies that human evaluation is not feasible to use for this study. SRG is measured by increasingly occluding the image and observing how the prediction changes. By occluding the pixels with the Least Influence First (LIF), the model prediction is expected to be similar until the influential pixels start getting occluded. Conversely, by occluding the pixels with the Most Influence First (MIF), the model prediction is expected to lower quickly. A good explanation should have a large area under the LIF prediction-occlusion curve and a small area under the MIF curve. LIF and MIF are equivalent to the insertion and deletion metrics used to evaluate RISE originally [9]. However, LIF and MIF have been shown to be inconsistent at ranking explanations as the occlusion type changes (e.g. random color, mean color, inpainting). The SRG metric, calculated as the difference between LIF and MIF, is more reliable [8]. This also eliminates some of the effects of occluded images being out-of-distribution. The connection between LIF, MIF, and SRG is visualized in Fig. 2.
Figure 2: Showcase of how LIF, MIF, and SRG metrics are calculated by steadily occluding the least or most influential pixels of an image and calculating the value of the top class predicted for the original image.
In this work, SRG is calculated by occluding the image over a total of 10 equal steps (from 0% occlusion in step 1 to 100% occlusion in step 10). The remaining pixels with the lowest or greatest attribution score for original top-class prediction are occluded in each step for LIF and MIF respectively. When there are many pixels with the same attribution, then pixels are chosen in an arbitrary deterministic order. Occlusion is performed by setting the pixels to the mean pixel value of the image, which mirrors one of the methods used for evaluating RISE originally [9]. The average of the original top-class prediction over these 10 images is then recorded as the LIF and MIF scores. The SRG score is calculated as $LIF-MIF$.
III Results and Analysis
The results consist of the LIF, MIF, and SRG metrics for every attribution pipeline. As this is too much data to present in this work, it is summarized as the average SRG metric for different parameter combinations. The tables show for some parameter combinations the SRG in % averaged over all pipelines with those combinations (i.e., the average result for all combinations of the unspecified parameters). The complete data is available online [^3] https://bit.ly/smooth-mask-perturbation.
The average SRG for pipelines with different sampling methods and sample sizes over the different attribution methods is shown in Table II. Unsurprisingly, increasing sample size yields improved performance. What is surprising is that random sampling significantly outperforms entropic and does so even for SHAP, for which entropic is the default sampling method. It is noteworthy that the attribution method (row-wise) in general has less impact on the SRG value than the sampling method (column-wise). Furthermore, some attribution methods seem to fail when used together with certain sampling methods and sample sizes. For non-random sampling with a sample size of 50 all attribution methods have the same performance. Under these circumstances, CIU, PDA, and RISE mathematically produce the same rankings of segments to which LIME and SHAP consistently converge.
TABLE II: The SRG in % for all pipelines with different combinations of sampling and attribution methods. The results are averaged for all pipelines with the specified parameters.
Attribution
Sampling Sample size CIU PDA LIME SHAP RISE
Only one $\leq 50$ 13.3 13.3 13.3 13.3 13.3
All but one $\leq 50$ 14.9 14.9 14.9 14.9 14.9
Random 50 - 16.0 6.8 6.8 16.0
Entropic 50 - 13.3 13.3 13.3 13.3
[…]
Entropic 400 - 9.0 15.6 17.3 15.0
[…]
Entropic 4000/8000 - 14.7 18.2 18.8 17.8
The results of different combinations of segmenting, perturbing, and attribution can be found in Table III. Again, the attribution calculation method has less impact on performance compared to the other parameters. Notably, average performance increases when attributing per-pixel for all combinations of segmenting, perturbing, and attribution methods. The Gaussian filter does not give a significantly different performance for grid segmentation compared to bilinear upsampling. However, as the Gaussian filter was implemented to mimic bilinear upsampling in this case, the results show that there is no harm in choosing one over the other. Instead, the benefit of smoothing with a Gaussian filter becomes obvious in the significant performance increase seen with SLIC over grid segmentation. As SLIC with pixel attribution is not possible with bilinear upsampling, this suggests that Gaussian filters enable the use of better segmentation methods without sacrificing performance on grid segmentation.
TABLE III: The SRG in % for all pipelines with different combinations of segmenting, perturbing, and attribution methods with either per-segment or per-pixel attribution. The results are averaged for all pipelines with the specified parameters.
Attribution
Segmenting + Perturbing Attribute per CIU PDA LIME SHAP RISE
Grid+bilinear Segment 11.7 14.9 14.9 14.5 15.9
Grid+bilinear Pixel 14.1 16.3 16.5 16.4 17.6
[…]
Grid+Gaussian Pixel 14.4 16.5 16.8 16.7 17.8
[…]
SLIC+Gaussian Pixel 16.8 17.6 18.2 18.3 18.8
CIU is not evaluated for random or entropic sampling, which have
greater average performance.
IV Discussion
Most works that introduce or compare perturbation-based image explanation methods do not examine the parameters of the explanation pipelines separately. This leads to a poor understanding of why a method performs well and can lead to questionable claims of some methods being better than others [28, 29]. The results of this work show that each parameter of the perturbation-based image explanations pipeline significantly impacts performance.
From the results, it is clear that using SLIC as opposed to grid segmenting improves performance significantly. Per-pixel attribution, which requires smooth masks, also improves performance compared to per-segment attribution. Finally, using SLIC with per-pixel attribution further improved performance. This shows that using Gaussian filters or other alternatives to bilinear smoothing is promising as it allows the use of better-performing segmentation methods. The choice of attribution calculation has a noticeably smaller impact on performance when compared to the choice of segmentation method or the difference between per-segment and per-pixel attribution. It is interesting that PDA and RISE attribution perform differently when they perform equally in the ideal case. It seems that, with a limited sample size, using the average decrease in prediction when a segment is excluded (PDA) is less stable than using the average increase when it is (RISE).
It is clear that an increased sample size does improve performance. However, the computation needed scales directly with the sample size. With ever-increasing computational demands by newer DNN models, even a low sample size let alone thousands of samples, may be unrealistic to presume for an explanation of a single decision. Developing perturbation-based methods that can give good explanations with low sample size is therefore a promising future direction. In some cases, such as medical diagnosis prediction, the need for and the value of explanations are likely high enough that it is worth increasing computational demands by factors of thousands.
The evaluation in this work relied on the explanation methods being separable into different parameters that could be combined in various ways. This is not always the case, even if the method otherwise produces sound explanations. For example, the original RISE implementation shifts the perturbation masks by some pixels so as not to center the same pixels every time. This approach works with the RISE attribution calculation since it can directly assign influence to pixels. However, this is not feasible for other methods, as such shifting has not been used in this evaluation. Additionally, the use of SRG and similar metrics requires attribution scores for individual features. For example, the CIU method can use intermediate concepts to explain combinations of features, which results in explanations for the combinations rather than the individual features.
This work shows how different parameters affect performance as measured by SRG. SRG measures how well the explanation reflects the actual impact of the attributed regions on the model. While ensuring that attribution maps explain the model correctly is helpful, it is not known whether it correlates with usefulness to humans. Existing evaluation on whether proxy metrics correspond to usefulness for humans is limited and requires further investigation [13]. It is also questionable whether proxy metrics can be used to compare different types of attribution maps at all. Attribution maps that indicate large regions, that provide heatmaps, or that show important edges and textures are likely to be interpreted differently by humans in a way not reflected by most proxy metrics.